OpenStudy (anonymous):

cos^2x=Sin^2x find x

OpenStudy (anonymous):

I know no solution, just made this up, can you help me with literature?

OpenStudy (anonymous):

1-sin^2x=sin^2x 1=2sin^2x 1/2=sin62x sinx=1/4

OpenStudy (anonymous):

sin^(-1) 1/4

OpenStudy (anonymous):

that's a rude way to attract people to literature qestion, sorry.

OpenStudy (anonymous):

\[\cos ^{2}x=\sin ^{2}x\] \[\cos ^{2}x-\sin ^{2}x=0\] \[\cos 2x=0=\cos \frac{ \pi }{ 2 }=\cos \left( 2n+1 \right)\frac{ \pi }{2 }\] \[2x=\left( 2n+1 \right)\frac{ \pi }{ 2 },x=\left( 2n+1 \right)\frac{ \pi }{4 },n=0,1,2,....\]

OpenStudy (anonymous):

you don't need all this. \[1-Sin^2x=Sin^2x~~~~~~~~~~~~~~~->~~~~~~~~~~~~1=2Sin^2x\]\[Sin ^{-1}(1/4)\]